Suppose that the total utility function of a consumer is given by TU(x,y) = 3x2 y and the prices of X and Y are 1 Birr and 2 Birr per unit, respectively. If the income of the consumer is 600 Birr and if he spends all of his income on the consumption of commodities of X and Y, find the optimum amount of X and Y that the consumer will consume at equilibrium and find MRTSx,y.


1
Expert's answer
2021-08-23T13:09:46-0400

Derive the budget constraint:

I = PxX + PyY

600 = X + 2Y

The utility maximizing rule is where ("\frac{MUx}{MUy}) = (\frac{Px}{Py})":

TU(x,y) = 3x2y


MUx = "\frac{\partial U} {\partial x} = 6xy"


MUy = "\frac{\partial U} {\partial y} = 3x^{2}"


"\frac{Px}{Py} = \frac{1}{2}"


"\frac{6xy}{3x^{2} } = \frac{1}{2}"


"\frac{2y}{x } = \frac{1}{2}"


Y = "\frac{x}{4}"


Substitute in the budget constraint:

600 = X + 2Y

600 = X + 2("\frac{x}{4}")

Multiply both sides by 4:

2400 = 4X + 2X

2400 = 6X

X = 400

Y = "\frac{x}{4}" = "\frac{400}{4}" = 100


TU(x,y) = (400,100)

The optimum amount of X and Y that the consumer will consume at equilibrium = 400 and 100

 

MRTSxy = "\frac{MUx}{MUy}"

MUx = 6xy

MUy = 3x2

MRTSxy = "\frac{6xy}{3x^{2} }" = "\frac{2y}{x }" = 2(100)/400 = "\frac{1}{2 }"


MRTSxy = "\frac{1}{2 }"